When 6.36 g of a solid mixture of (NH4)2CrO7 and (NH4)2CO3 was subjected to decomposition by heat, the mass loss was 4.84g.
Calculate the no. of moles of each component in the initial mixture.
Calculate the total mass of h2o evolved.
Please help me with this question in steps
1
Expert's answer
2020-05-01T14:22:00-0400

Let the number of moles of (NH4)2Cr2O7 be x, and that of (NH4)2CO3 be y. Multiplying by molar masses of the compounds, we have the first equation for the total mass of solid mixture:

6.36 = 252*x + 96*y

Now write the balanced reaction equations:

(NH4)2Cr2O7(s) = Cr2O3 (s) + N2 (g) + 4  H2O (g)

(NH4)2CO3(s) = 2 NH3(g) + H2O(g) + CO2(g)

Mass loss is due to gases evolution. The number of moles of gases evolved are: N2 = x;

H2O = 4x+y; NH3 = 2y; CO2 = y. Multiplying by molar masses of each gas, we get the second equation for the mass loss:

4.84 = 28*x + 18*(4x+y) + 17*2y + 44*y = 100*x + 96*y

Solving the two equations, we find x and y:

x = 0.01 mole of (NH4)2Cr2O7,

y = 0.04 mole of (NH4)2CO3.

The total mass of H2O evolved is 18*(4*0.01+0.04) = 1.44 g